OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. y – 4 = 2(x + 3) y + 4 = 2(x – 3) y + 3 = 2(x – 4) y – 3 = 2(x + 4)

OpenStudy (anonymous):

Hint: Point-Slope form. y - y1 = m(x - x1) (x1, y1) m = slope

OpenStudy (anonymous):

so y – 3 = 2(x + 4) @ jazy

OpenStudy (anonymous):

Not quite. What are your x and y values in the given point?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

That is your y value: (x, y) (-3, 4) Okay, so in point-slope form: y - y1 = m(x - 1) Substitute -3 and 4 into the corresponding x1 and y2.

OpenStudy (anonymous):

What do you get? Also, remember that two negatives(-) equal a positive(+). (-)(-) = (+)

OpenStudy (anonymous):

y + 3 = 2(x – 4)

OpenStudy (anonymous):

waitt

OpenStudy (anonymous):

-3 is the x-value

OpenStudy (anonymous):

y – 4 = 2(x + 3)

OpenStudy (anonymous):

Right! (:

OpenStudy (anonymous):

thanks.. can u help me some more

OpenStudy (anonymous):

I'll try to help, I have to go soon though.